OpenStudy (anonymous):

What is the sum of a 23–term arithmetic sequence where the first term is 9 and the last term is 163?

OpenStudy (amistre64):

use the formula:\[\frac{n(first+last)}{2}\]

OpenStudy (anonymous):

so (9+163)

OpenStudy (anonymous):

/2

OpenStudy (amistre64):

good, times 23, since thats the number of terms

OpenStudy (amistre64):

way back when, before they had notations for this, they used words; and they had 2 different thrms for it that amounted to if the number of terms was even or odd even number:\[\frac{n}{2}(first+last)\] if it was odd\[n\frac{first+last}{2}\]